if the population consists of number 1,2,3,4,5 and 6,find the mean of the samples of size 3.construct a sampling distribution


1
Expert's answer
2022-03-25T15:20:03-0400

Mean of the samples




Sampling distribution

"P(X=2)=P(X=2.33)=P(X=5)=P(X=4.67)=\frac{1}{20}"

"P(X=2.67)=P(X=4.33)=\frac{2}{20}"

"P(X=3)=P(X=3.33)=P(X=3.67)=P(X=4)=\frac{3}{20}"